An airplane is preparing to land at an airport. It is 52200 feet above the ground and is descending at the rate of 3300 feet per minute. At the same​ airport, another airplane is taking off and will ascend at the rate of 2500 feet per minute. When will the two airplanes be at the same altitude and what will that altitude​ be?
Answer by ankor@dixie-net.com(22740)

An airplane is preparing to land at an airport. It is 52200 feet above the ground and is descending at the rate of 3300 feet per minute.
At the same​ airport, another airplane is taking off and will ascend at the rate of 2500 feet per minute.
When will the two airplanes be at the same altitude and what will that altitude​ be?
:
let t = no of minutes when they are at the same altitude
:
52200 - 3300t = 2500t
52200 = 3300t + 2500t
52200 = 5800t
t = 52200/5800
t = 9 minutes they will be at the same altitude
:
The altitude:
9*2500 = 22500 ft
:
Check on other train
52200 - 9(3300) =
52200 - 29700 = 22500 ft